# Security Weaknesses of Copilot Generated Code in GitHub

### v1

@@ 4. Results
We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then provide a detailed presentation of the final results for each RQ.
@@ 4.1. RQ1: How secure is the code generated by Copilot?
Approach. To answer this RQ, we collected 435 code snippets generated by Copilot from GitHub projects. These snippets cover six common programming languages. We used two static analysis tools (CodeQL + another language-dedicated tool) to scan and analyze the code snippets and then combine the results obtained from the two tools. The aim is to achieve a better coverage of security issues. Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
In the analysis results obtained from the CodeQL tool, three types of warnings were used to describe the detected weaknesses: Recommendation, which provides suggestions for improving code quality; Warning, which alerts to potential weaknesses that could cause code to run abnormally or unsafely; and Error, which is the highest level of warning and alert to inform that the error could cause code to fail to compile or run incorrectly. Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored the other code quality recommendations. For the scanning results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot generated code snippets based on the comment that appears before the method. We provide an example of the scan results filtration in Fig. 6. In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start and end lines of the scan results that suggested a security issue. In Step 2, we located the code at Line 103 and found no comment indicating that Copilot generated it. In Step 3, we also found that the code snippet generated by Copilot in the file was located at Lines 226 to 239, and we then determined that the code snippet generated by Copilot did not cause this security issue and discarded this scan result from further analysis. Finally, we aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
@@
Results. Table 5 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses. From the statistical results, we found that out of the 435 code snippets generated by Copilot, 35.8% of them have security weaknesses, regardless of the programming language. There is a higher proportion of security weaknesses in Python and JavaScript code, which are the most popular Copilot languages developers use (Stackscale, 2021). Out of the 251 Python code snippets, 39.4% have security weaknesses. Among the 79 JavaScript code snippets we collected, 27.8% have security weaknesses. Among all programming languages, C++ code snippets have the highest proportion of security weaknesses, reaching 46.1%. Go also has a relatively high proportion of security weaknesses, at 45.0%. In comparison, the proportion of files with security issues is lower for C# and Java code, at 25% and 23.2%, respectively.
@@
Python 251 99 39.4%
JavaScript 79 22 27.8%
Java 43 10 23.2%
C++ 26 12 46.1%
Go 20 9 45.0%
C# 16 4 25%
Total 435 156 35.8%
@@ 4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
Approach. To answer RQ2, we processed the results of the scans conducted for RQ1, eliminating duplicate security issues detected at the same code snippet location. In total, we identified 600 security weaknesses across 435 code snippets. Table 6 shows the number of security weaknesses found in code files of different programming languages.
@@
Python 99 352
JavaScript 22 93
Java 10 56
C++ 12 70
Go 9 18
C# 4 11
Total 156 600
@@
For each code snippet, we used CWEs to classify the security issues identified by the static analysis. Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, [n. d.]). Some static analysis tools we used, such as Bandit and Gosec, provide a CWE ID corresponding to the detected security issues in their scan results. For other scan tools that do not directly give a CWE ID, such as Codeql, ESLint, and FindBugs, we manually associated the provided security issue information with a CWE ID, which is detailed below. The QL queries used in CodeQL often point to a specific CWE, and the scanning process typically displays the CWE-ID associated with the QL query. Although only the name of the QL query is shown in the scan results, we can manually correlate the name field with the CWE ID. For example, the result shows “Hard-coded credential in API call”, and we know that Hardcodedcredential.ql query belongs to CWE-798, this security issue can be mapped to the same CWE (798). In addition, while FindBugs and ESLint also identify security issues, their scan results only provide descriptions of the security issues. We manually associated the descriptions of the security issues in the results with relevant CWE descriptions to determine the specific CWE category to which these security issues belong. Initially, two authors independently matched each description of the security issue with a CWE ID. In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment. This process continued until all the descriptions of the security issues in the results were matched with CWE IDs. Table 7 shows the list of manually matched CWE IDs and warning messages from CodeQL, ESLint, and FindBugs results. In the final stage, we performed a statistical analysis of CWE weaknesses in 156 code snippets that contained security weaknesses.
@@
Results. Table 8 shows the distribution of CWEs in the code snippets, including the number of code snippets that contain a certain CWE (Related Snippets) and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”). Note that one related code snippets may contain multiple instances of specific CWE. In total, we found 600 CWEs in 435 code snippets. These security weaknesses were related to 42 types of CWE, indicating that developers face a variety of security weaknesses when using Copilot. CWE-78: OS Command Injection is the most frequently occurred CWE, as it was detected in 15 code snippets (representing 14% of the security weaknesses), followed by CWE-330: Use of Insufficiently Random Values, CWE-703: Improper Check or Handling of Exceptional Conditions , CWE-400: Uncontrolled Resource Consumption and CWE-502: Deserialization of Untrusted Data. Some CWEs appeared less frequently, such as CWE-95: Eval Injection, and CWE-22: Improper Limitation of a Pathname to a Restricted Directory.
Additionally, many CWEs occur with a probability of less than 1%, for example, CWE-176: Improper Handling of Unicode Encoding, CWE-312: Cleartext Storage of Sensitive Information, and CWE-326: Inadequate Encryption Strength. This indicates that the types of security issues are closely related to the specific scenarios in which developers use Copilot and make the security issues become apparent, emphasizing the importance of maintaining vigilance and caution when programming.
Table 7. The warning messages of the scan results by CodeQL, ESLint, and FindBugs manually matched with corresponding CWE IDs
Tool Warning Message CWE-ID
CodeQL
Reflected server-side cross-site scripting
CWE-79
CodeQL
Flask app is run in debug mode
CWE-215
CodeQL
Clear-text logging of sensitive information
CWE-532
CodeQL
Clear-text storage of sensitive information
CWE-312
CodeQL
Information exposure through an exception
CWE-209
CodeQL
Request without certificate validation
CWE-295
CodeQL
Assignment to constant
CWE-682
CodeQL
Log injection
CWE-117
CodeQL
Identical operands
CWE-570
CodeQL
Incomplete string escaping or encoding
CWE-176
CodeQL
DOM text reinterpreted as HTML
CWE-79
CodeQL
Arbitrary file write during archive extraction (“Zip Slip”)
CWE-22
CodeQL
Hard-coded credential in API call
CWE-798
CodeQL
Hard-coded credentials
CWE-798
CodeQL
Uncontrolled data used in path expression
CWE-22
CodeQL
Resource not released in destructor
CWE-416
CodeQL
Missing Dispose call on local IDisposable
CWE-690
CodeQL
Use of the return value of a procedure
CWE-252
CodeQL
Dereferenced variable may be null
CWE-476
ESLint
Generic Object Injection Sink
CWE-502
ESLint
Function Call Object Injection Sink
CWE-20
ESLint
Unsafe Regular Expression
CWE-20
ESLint
Variable Assigned to Object Injection Sink
CWE-95
FindBugs
M S Dm: Hardcoded constant database password
CWE-798
FindBugs
H I Dm: Found reliance on default encoding
CWE-116
FindBugs
M D ICAST: Integral division result
CWE-682
FindBugs
H C IL: There is an apparent infinite loop
CWE-835
FindBugs
M B RV:Exceptional return value of java.io.File.mkdirs()ignored
CWE-252
FindBugs
M D NP: Possible null pointer dereference
CWE-476
FindBugs
M B ODR: DatabaseOperate.readDatabase() may fail to close Connection
CWE-404
@@ H C IL: There is an apparent infinite loop
CWE-ID # Related Code Snippets Frequency of Specific CWE Percentage
CWE-78
15
84
14.0%
CWE-330
34
81
13.5%
CWE-703
20
78
13.0%
CWE-398
11
60
10.0%
CWE-502
15
56
9.3%
CWE-400
22
50
8.3%
CWE-20
10
19
3.1%
CWE-252
2
13
2.1%
CWE-259
6
13
2.1%
CWE-404
3
13
2.1%
CWE-451
3
13
2.1%
CWE-682
2
13
2.1%
CWE-116
3
11
1.8%
CWE-690
3
9
1.5%
CWE-798
5
9
1.5%
CWE-561
4
8
1.3%
CWE-95
4
8
1.3%
CWE-22
4
6
1.0%
CWE-327
5
5
¡1%
CWE-532
3
5
¡1%
CWE-563
4
4
¡1%
CWE-605
4
4
¡1%
CWE-89
3
4
¡1%
CWE-295
1
3
¡1%
CWE-476
1
3
¡1%
CWE-775
2
3
¡1%
CWE-117
1
2
¡1%
CWE-209
1
2
¡1%
CWE-215
2
2
¡1%
CWE-416
1
2
¡1%
CWE-570
2
2
¡1%
CWE-664
1
2
¡1%
CWE-676
2
2
¡1%
CWE-79
1
2
¡1%
CWE-94
2
2
¡1%
Others
=1
¡1%
42 Types Total: 600
@@ 4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
Approach. The code in our collected dataset was generated between June 2021 and June 2023. To compare whether the security issues in Copilot generated code are widespread in this period, we chose MITRE 2022 CWE Top-25 list (The MITRE Corporation, 2022) as our baseline. Then, we compared the CWEs obtained in RQ2 with the CWE Top-25.
Results. The distribution of CWEs found compared to the MITER list is shown in Table 9. The results show that the CWE weaknesses present in the code generated by Copilot belong to eleven CWE types included in the MITER CWE Top-25 list. This means these are present issues and are currently among the most common and serious security weaknesses in practice. It is worth noting that the 237 security issues present in the code snippet correspond to these 11 CWEs, while another 31 CWEs cover the remaining 363 security issues. This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot. Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their codebase. At the same time, we can see that CWE-78: OS Command Injection is the most frequently occurring weakness from the Top-25 security weaknesses, ranking sixth in the Top-25 list and first in our RQ2 results. Although CWE-400: Uncontrolled Resource Consumption is ranked towards the end of the Top-25 list, it is one of the weaknesses with a high occurrence frequency. Some CWEs with a higher ranking in the Top-25 list do not appear frequently in Copilot generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
Table 9. The CWEs that belong to the 2022 CWE Top-25 list
CWE-ID Description # Related Snippets Frequency
CWE-78
OS Command Injection
15
84
CWE-502
Deserialization of Untrusted Data
15
56
CWE-400
Uncontrolled Resource Consumption
22
50
CWE-20
Improper Input Validation
10
19
CWE-798
Use of Hard-coded Credentials
5
9
CWE-22
Improper Limitation of a Pathname to a Restricted Directory
4
6
CWE-89
SQL Injection
3
4
CWE-476
NULL Pointer Dereference
1
3
CWE-94
Code Injection
2
2
CWE-416
Use After Free
1
2
CWE-79
Cross-site Scripting
1
2
Total 237
@@ 5.1. Interpretation of Results
Among the 435 code snippets generated by Copilot, we found that 35.8% of these code snippets contain security weaknesses. Those weaknesses appear in all six top-used programming languages supported by Copilot. Furthermore, when it comes to the occurrence of security issues in code snippets of different programming languages, it is important to analyze them in conjunction with the popularity of the languages (mend.io, 2023). In code snippets written in languages like Python and JavaScript, which are frequently used with Copilot, there may be a slightly higher number of security issues. However, overall the proportion of security issues across these six languages ranges from 25% to 45%, showing no significant difference.
Besides, we also found that CWE-502: Deserialization of Untrusted Data and CWE-400: Uncontrolled Resource Consumption problems mainly appeared in code snippets written in Python and JavaScript. This could be attributed to certain features that made their code more flexible, such as dynamic typing and dynamic interpretation. Therefore, developers should pay special attention to the security of their JavaScript and Python-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks. The results of RQ1 suggest that in practical production, although Copilot can help developers write code faster and increase productivity, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
@@
After conducting a security evaluation of 425 code snippets generated by Copilot, a total of 600 security weaknesses were identified, involving 42 CWE types, which is around 10% of the CWEs (439 CWE types) in software development (The MITRE Corporation, 2023). This may be due to the reason that Copilot generates code in different programming languages and application scenarios, and a wide variety of application scenarios may lead to various types of security issues. In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot. This set of 42 CWE types covers many types of security issues, and Table 10 shows the types of security issues that these 42 CWEs are relevant to.
Table 10. The CWEs and Types of Security issues
@@
Web security issue
CWE-79, CWE-94, CWE-690, CWE-732
Access control issue
CWE-252, CWE-259, CWE-327, CWE-338
Input validation and representation issue
CWE-20, CWE-89, CWE-116
Command injection issue
CWE-78, CWE-563, CWE-835
SQL injection issue
CWE-89, CWE-95
File handling issue
CWE-22, CWE-570
Insecure storage
CWE-502, CWE-775
Improper error handling
CWE-398, CWE-400
Encryption issue
CWE-327
Memory management issue
CWE-416
Buffer errors
CWE-476
Insecure random number
CWE-330
Incorrect type conversion
CWE-703
The diversity of security weaknesses indicates that developers using Copilot face various security risks. These risks are diverse, covering different development environments and application scenarios. At the same time, it also reflects the inevitability of security weaknesses in Copilot generated code. Developers need to have corresponding security awareness and skills and take appropriate security measures to avoid these risks in a timely and targeted manner. In addition, we can see that developers often encounter CWE-78: OS Command Injection, CWE-330: Use of Insufficiently Random Values, andCWE-703: Improper Check or Handling of Exceptional Conditions, CWE-502: Deserialization of Untrusted Data, and CWE-400:Uncontrolled Resource Consumption, which appear in multiple code snippets and have a high frequency of occurrence. This can remind developers to take timely and targeted security measures to mitigate these risks. For example, developers should perform adequate validation of user inputs. In addition to this, it is also necessary to restrict the program’s permissions so that they only access essential resources. The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers be aware of security aspects of code generated by Copilot and take appropriate measures to address the security weaknesses in an informed manner.
@@
As shown in Table 9, eleven of the CWEs in Copilot generated code can be found in the 2022 CWE Top-25 list, covering more than 237 security issues (39.5% of 600 identfied CWEs) in our dataset. This indicates that the commonly acknowledged top 25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot. Therefore, developers need to pay special attention to these frequently occurring weaknesses and take corresponding measures to avoid and fix them. We also observed that some vulnerabilities from the CWE top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers. GitHub is gradually enhancing the security of Copilot and its underlying model (Codex) (GitHub, 2023). We also identified 31 security weaknesses in the code that do not belong to the CWE Top-25 list. Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them. For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset. This security weakness is not commonly found in code and only occurs when specific users have certain permissions. However, it can lead to significant security risks when it does occur. Developers should also be aware of these less common security weaknesses to fully protect their code from attacks.

### v2

@@ 4. Results
We present the results of three RQs formulated in Section 3.1 below. For each RQ, we first explain how we analyzed the collected code snippets to answer the RQ. We then answer each of our three RQs.
@@ 4.1. RQ1: How secure is the code generated by Copilot?
Approach. To answer this RQ, we collected 452 Python and JavaScript code snippets generated by Copilot from GitHub projects. We used two static analysis tools (CodeQL + another language-dedicated tool, Bandit for Python and ESLint for JavaScript) to scan and analyze the code snippets and then combine the results obtained from the two tools. The aim is to achieve better coverage of security issues. Therefore, as long as one of the tools detected the presence of a security issue, the code snippet was considered vulnerable.
We identified three types of warnings from CodeQL analysis:
  • Recommendation, which provides suggestions for improving code quality;
  • Warning, which alerts to potential weaknesses that could cause code to run abnormally or unsafely;
  • Error, which is the highest level of warning and alert to inform that the error could cause code to fail to compile or run incorrectly.
Since our research primarily focused on security weaknesses, we only counted code snippets that had warnings and errors, and we ignored recommendations on code quality. After obtaining the initial scan results related to security issues, we manually checked the scan results. For the scan results from the Repository label, we marked them as 1 (security issue exists) and 0 (no security issue exists). For the scan results from Code label, we marked them as 1 (security issue exists) and 0 (no security issue exists or the code snippet with a security issue was not generated by Copilot). Figure 7 shows the outcomes of the manual verification on the initial security scan results under the Repository and Code labels.
When manually checking the scan results from the Code label, we also needed to identify whether the security issues obtained from the scan were from Copilot-generated code snippets based on the comment (prompt message) that appears before the method. We provide a working example of the filtration of the scan results in Figure 8. In Step 1, we first went to the corresponding file to locate the specific code snippet based on the start line numbers of the scan results. In Step 2, we located the code at Line 149 in aggregation.py. We found that this code does have a security issue and determined that it was generated by Copilot based on the prompt messages above. Consequently, we marked the corresponding security scan result as “1”. We also located the code at Line 145 in _lytools.py. We found that this code had a security issue, but there were no prompt messages that would indicate that it was generated by Copilot. Consequently, we marked the corresponding security scan result as “0” and discarded this scan result from further analysis.
Finally, we kept the results marked as 1 and aggregated the filtered results obtained using multiple analysis tools to calculate the number of code snippets with security issues detected.
(a) Results from the Repository label
(b) Results from the Code label
Figure 7. Manual verification of the initial security scan results for the Repository and Code label
@@
Results. Table 3 shows the numbers of code snippets for different types and the numbers and percentages of code snippets with security weaknesses. From the statistical results, we found that out of the 452 code snippets generated by Copilot, 25.9% of them have security weaknesses. Among the Python and JavaScript code snippets, Python exhibits a higher proportion of security issues compared to JavaScript. Out of the 277 Python code snippets, 91 (32.8%) have security weaknesses. Among the 175 JavaScript code snippets we collected, 43 (24.6%) have security weaknesses. Note that one related code snippet may contain multiple instances of a specific CWE.
@@
Python 277 91 32.8%
JavaScript 175 43 24.6%
Total 452 134 29.6%
@@ 4.2. RQ2: What security weaknesses are present in the code snippets generated by Copilot?
Approach. To answer RQ2, we processed the scan results collected by RQ1 that were manually checked to contain vulnerabilities. In total, we found 544 security issues in 452 code snippets, and Table 4 shows the number of security weaknesses found in code files of different programming languages.
@@
Python 91 277
JavaScript 43 175
Total 134 544
@@
For each code snippet, we used CWEs to classify the security issues identified by the static analysis tools. Each CWE has a unique ID and a set of related descriptions, including its potential impact and how to detect and fix the CWE (The MITRE Corporation, 2023b). To more accurately identify the type of security issue with the corresponding CWE, we did not directly accept the provided information of the static analysis tools (such as the CWE number reported by Bandit). Instead, we manually identified the type of security issue at the specific location reported in the warning message of the static analysis tool and matched it with the corresponding CWE number. Initially, two authors independently matched each description of the security issue with a CWE ID. In case of disagreement, a discussion was initiated between the two authors, and one other author (a security expert) was then involved to provide his assessment. This process continued until all the descriptions of the security issues in the results were matched with CWE IDs. In the final stage, we performed a statistical analysis of CWE weaknesses in 134 code snippets that contained security weaknesses.
@@
Results. Table LABEL:Distribution_of_CWEs_in_code_snippets shows the distribution of CWEs in the code snippets and the total number of occurrences (Frequency) of the CWE in the code snippets (we put those CWEs whose Frequency = 1 in “Others”). In total, we found 544 CWEs in 452 code snippets. These security weaknesses were related to 38 CWEs, indicating that developers face a variety of security weaknesses when using Copilot. CWE-330: Use of Insufficiently Random Values is the most frequently occurring CWE, as it represents 23.3% of the security weaknesses, followed by CWE-94: Improper Control of Generation of Code (‘Code Injection’) , CWE-78: OS Command Injection and CWE-95: Improper Neutralization of Directives in Dynamically Evaluated Code (‘Eval Injection’). Some CWEs appear less frequently, for example CWE-117: Improper Output Neutralization for Logs only occurred twice. Furthermore, many CWEs occur with a frequency of less than 1%, for example, CWE-22: Improper Limitation of a Pathname to a Restricted Directoryand CWE-770: Allocation of Resources Without Limits or Throttling. This indicates that the types of security issue are closely related to the specific scenarios in which developers use Copilot, emphasizing the importance of maintaining vigilance and caution when programming. Besides, Table 6 presents the top 5 CWEs that appear in Python and JavaScript code snippets.
@@
CWE-ID Frequency of Specific CWE Percentage
CWE-330 127 23.3%
CWE-94 115 21.1%
CWE-78 81 14.9%
CWE-95 30 5.5%
CWE-772 23 4.2%
CWE-89 15 2.8%
CWE-457 15 2.8%
CWE-259 14 2.6%
CWE-670 13 2.4%
CWE-396 11 2.0%
CWE-617 8 1.5%
CWE-628 7 1.3%
CWE-563 7 1.3%
CWE-561 7 1.3%
CWE-798 6 1.1%
CWE-703 6 1.1%
CWE-502 6 1.1%
CWE-20 6 1.1%
CWE-312 5 0.9%
CWE-605 4 0.7%
CWE-295 4 0.7%
CWE-252 4 0.7%
CWE-367 3 0.5%
CWE-327 3 0.5%
CWE-185 3 0.5%
CWE-79 2 0.4%
CWE-682 2 0.4%
CWE-665 2 0.4%
CWE-400 2 0.4%
CWE-215 2 0.4%
CWE-200 2 0.4%
CWE-117 2 0.4%
CWE-116 2 0.4%
Others =1 ¡1%
38 Types Total: 544
Table 6. Top 5 CWEs in Python and JavaScript
Rank CWE Type
Python JavaScript
1 CWE-330 (Use of Insufficiently Random Values Weakness) CWE-94 (Improper Control of Generation of Code)
2 CWE-78 (Improper Neutralization of Special Elements used in an OS Command) CWE-95 (Improper Neutralization of Directives in Dynamically Evaluated Code)
3 CWE-772 (Missing Release of Resource after Effective Lifetime) CWE-563 (Assignment to Variable without Use)
4 CWE-89 (Improper Neutralization of Special Elements used in an SQL Command) CWE-20 (Improper Input Validation)
5 CWE-259 (Use of Hard-coded Password) CWE-185 (Incorrect Regular Expression)
@@ 4.3. RQ3: How many security weaknesses belong to the CWE Top-25?
Approach. The code snippets in our collected dataset were mainly generated in 2022 and 2023. To compare whether the security issues in Copilot-generated code are widespread in this period, we chosed MITRE 2023 CWE Top-25 list (The MITRE Corporation, 2023a) as our baseline. In addition, MITRE maintains another list of “Stubborn Weaknesses in the CWE Top 25” (The MITRE Corporation, 2023d), which includes those CWEs that have consistently appeared in the Top 25 most dangerous software weaknesses of the past five years (2019-2023). We compared the CWEs obtained in RQ2 with the 2023 CWE Top 25 and “Stubborn Weaknesses” in the CWE Top 25.
Results. The distribution of CWEs found compared to the MITER list is shown in Table 7. The results show that the CWE weaknesses present in the code generated by Copilot belong to eight CWEs. We highlight those CWEs that are “Stubborn Weaknesses” (SW) in bold. This means these are present issues and are currently among the most common and serious security weaknesses in practice. It is worth noting that the 218 security issues present in the code snippet correspond to these eight CWEs, while another 30 CWEs cover the remaining 326 security issues. This indicates that the CWE Top-25 weaknesses are also prevalent in the code generated by Copilot. At the same time, we can see that CWE-78: OS Command Injection occurs more frequently in Copilot-generated code. CWE-94: Improper Control of Generation of Code is in the list of the CWE Top-25 (The MITRE Corporation, 2023a), and meanwhile it is also one of the weaknesses with a high occurrence frequency in our results. Some CWEs with a higher ranking in the Top-25 list do not frequently appear in Copilot-generated code, such as CWE-79: Cross-site Scripting and CWE-89: SQL Injection.
Table 7. The CWEs that belong to the 2023 CWE Top-25 list
CWE-ID Description Is SW # Related Snippets Frequency
CWE-94 Improper Control of Generation of Code No 34 115
CWE-78 Improper Neutralization of Special Elements used in an OS Command Yes 18 81
CWE-798 Use of Hard-coded Credentials Yes 3 6
CWE-502 Deserialization of Untrusted Data Yes 2 6
CWE-20 Improper Input Validation Yes 3 6
CWE-79 Improper Neutralization of Input During Web Page Generation Yes 2 2
CWE-276 Incorrect Default Permissions No 1 1
CWE-22 Improper Limitation of a Pathname to a Restricted Directory Yes 1 1
Total 218
@@ 5.1. Interpretation of Results
Among the 452 code snippets generated by Copilot, we found that 29.6% of these code snippets contain security weaknesses. Although in our results, the proportion of security issues in Python code snippets is higher than in JavaScript, there is no significant difference. Using Copilot to write code in Python or JavaScript can generally lead to security issues. This could be attributed to features that made their code more flexible, such as dynamic typing and interpretation. From a Copilot point of view, the generated code does not need to reason about the whole program because it is dynamically typed, consequently, it does not need to read the flow of the whole program to generate working suggestions. In summary, developers should pay special attention to the security of Copilot-generated code, taking appropriate measures to validate input data and manage resources effectively to minimize security risks. The results of RQ1 suggest that in practical production, Copilot can help developers write code faster and increase productivity (GitHub, 2024; Perry et al., 2022). Still, additional security assessments and fixes are also required to ensure that the generated code does not introduce potential security risks.
@@
After conducting a security evaluation of 452 code snippets generated by Copilot, a total of 544 security weaknesses were identified, involving 38 CWEs, which is around 10% of the CWEs (439 CWEs) (The MITRE Corporation, 2023c). This may be because Copilot generates code in different programming languages and application scenarios. In addition, since the Copilot base model (Codex) is trained on publicly available data that potentially contain various types of security weaknesses, this can lead to the presence of multiple CWEs in the generated code by Copilot. This set of 38 CWEs covers many security issues, and Table 8 shows the types of security issues to which these CWEs are relevant.
Table 8. CWEs and security issue types in Copilot-generated code
@@
Data Neutralization Issues CWE-94, CWE-78, CWE-117, CWE-89
Resource Management Errors CWE-772, CWE-502
Error Conditions, Return Values CWE-396, CWE-617
Bad Coding Practices CWE-628, CWe-563
Credentials Management Errors CWE-798
Information Management Errors CWE-312, CWE-215
Authentication Errors CWE-295
Concurrency Issues CWE-367
File Handling Issues CWE-22
The diversity of security weaknesses indicates that developers using Copilot face various security risks. These risks are diverse, covering different development environments and application scenarios. In addition, we find that CWEs in Python are mainly related to data processing and system calls. In contrast, CWEs in JavaScript are often associated with dynamic code generation problems and security issues in web development. This may be because the two languages are designed differently. JavaScript is more commonly used for web development, while Python is more widely used in areas such as data processing and scientific computing. Overall, while Python and JavaScript differ in some common types of security weaknesses, they require developers to be aware of and take timely and targeted security measures to mitigate these risks. For example, developers should perform adequate validation of user inputs. It is also necessary to restrict the program’s permissions so that they only access essential resources.
The results of RQ2 reveal the security weaknesses that developers may encounter in an actual production environment and their frequency of occurrence, which can help developers to be aware of the security aspects of the code generated by Copilot and to take appropriate measures to address the security weaknesses in an informed manner.
@@
As shown in Table 7, eight of the CWEs in the Copilot-generated code can be found in the 2023 CWE Top-25 list, covering more than 218 security issues (40.1% of 544 identified CWEs) in our dataset. This indicates that the commonly acknowledged CWE Top-25 weaknesses in software development, which are considered the most prevalent and dangerous, are also prevalent in the code generated by Copilot. Therefore, developers using Copilot must pay close attention to these weaknesses and take appropriate measures to prevent them before they are integrated into their code. Meanwhile, we found that some higher-ranked CWEs in the CWE Top-25 list, such as CWE-89: SQL Injection, did not frequently appear in our scan results. This could be because the generated code does not establish connections with external resources (e.g., databases). Besides, we also observed that some vulnerabilities from the CWE Top-25 list were not detected in our scans, indicating that Copilot may sanitize and prevent specific weaknesses from being suggested to developers. GitHub is gradually improving the security of Copilot and its underlying model (Codex) (GitHub, 2023). However, it has also been found that Copilot’s security layer can potentially handle some CWEs better than others, leaving applications vulnerable to some critical vulnerabilities (Majdinasab et al., 2024).
We also identified 30 security weaknesses in the code that do not belong to the CWE Top-25 list. Although these less common security weaknesses may not be as widespread as CWE Top-25, attackers can still exploit them. For example, we only detected one instance of CWE-732: Incorrect Permission Assignment for Critical Resource in our dataset. This security weakness is not commonly found in code and only occurs when specific users have certain permissions. However, it can lead to significant security risks when it does occur. Developers should also be aware of these less common security weaknesses to fully protect their code from attacks.